OpenStudy (anonymous):

A scientist suggests keeping indoor air relatively clean as follows: provide 2 or 3 pots of flowers for every 100 square feet of floor space under a ceiling of 8 ft. If your classroom has an 8 ft ceiling and measures 35 ft by 40 ft, how many pots should it have?

OpenStudy (anonymous):

I believe this is the correct way to work on this problem. 10x10=100sqft 2 or 3 pots 34x40=1360sqft pots=p 2/100 = p/1360 100p =2720 p = 2720/100 Then Simplify 3/100 = p/1360 100p = 4080 p = 4080/100 Then Simplify Did that help?

OpenStudy (anonymous):

no\